OpenStudy (ohohaye):

Please help me Show that the point (7/25,24/25) is on the unit circle

OpenStudy (ohohaye):

So just a^2+b^2=c^2?

OpenStudy (janu16):

yea

OpenStudy (ohohaye):

Oh wait, I think you got the wrong problem. This is the point you have to find on the unit circle \[(\frac{ 7 }{ 25 },\frac{ 24 }{ 25 })\]

OpenStudy (ohohaye):

@Janu16

OpenStudy (janu16):

no i got it right. but heres the other way The center of the unit circle is (0,0). If the distance between the center and the given point turns out to be 1, then the point is on the unit circle. The distance of the given point from (0,0) in this case works out to be \[\sqrt{\left(\begin{matrix}7 \\ 25\end{matrix}\right)^2+\left(\begin{matrix}24 \\ 25\end{matrix}\right)^2}\] \[=\sqrt{49+576}\div625\] =1

OpenStudy (janu16):

@ohohaye

OpenStudy (ohohaye):

oh, I'm sorry, I've never seen someone solve it how you were solving it in the first part

OpenStudy (janu16):

but didyou get 2nd part?

OpenStudy (janu16):

@ohohaye

OpenStudy (ohohaye):

I'm confused as to this part

OpenStudy (ohohaye):

@Janu16

OpenStudy (janu16):

so 49+576 is 625 and when you divide 625 by 625 its 1

OpenStudy (janu16):

@ohohaye

OpenStudy (ohohaye):

Ok, thank you